working together, matt and juan can restock a store's shelves in 2 hours 24 minutes. alone juan needs two hours longer than matt needs. how long does matt need to restock the shelves when working alone
Answer by checkley77(12844)

2.4=x(x+2)/(x+x+2)
2.4=x^2+2x/2x+2
x^2+2x=2.4(2x+2)
x^2+2x=4.8x+4.8
x^2+2x-4.8x-4.8=0
x^2-2.8x-4.8=0
(x-4)(x+1.8)=0
x-4=0
x=4 ans. for Matt working alone.
4+2=6 ans. for Juan working alone.
Proof:
2.4=4*6/(4+6)
2.4=24/10
2.4=2.4